OpenStudy (thecatman):

The formula K equals C plus two hundred seventy-three shows how to estimate the temperature K measured in degrees Kelvin (°K) when you know the temperature C measured in degrees Celsius (°C). If room temperature is 295°K, what is room temperature in °C? -22°C 22°C 568°C -568°C

OpenStudy (thecatman):

@jennaknight03

OpenStudy (whpalmer4):

K = C + 273 is your equation. You know the temperature is 295 K and want to find the value for C, so plug in what you know and solve for what you don't. 295 = C + 273 C=?

OpenStudy (thecatman):

22

OpenStudy (anonymous):

@TheCatMan you need help you never need help you are the one giving the help

OpenStudy (anonymous):

it is awesome you are getting help